# Mechanistic Interpretability for AI Safety A Review

### v2

@@ Abstract
Understanding AI systems’ inner workings is critical for ensuring value alignment and safety. This review explores mechanistic interpretability: reverse engineering the computational mechanisms and representations learned by neural networks into human-understandable algorithms and concepts to provide a granular, causal understanding. We establish foundational concepts such as features encoding knowledge within neural activations and hypotheses about their representation and computation. We survey methodologies for causally dissecting model behaviors and assess the relevance of mechanistic interpretability to AI safety. We examine benefits in understanding, control, and alignment, along with risks such as capability gains and dual-use concerns. We investigate challenges surrounding scalability, automation, and comprehensive interpretation. We advocate for clarifying concepts, setting standards, and scaling techniques to handle complex models and behaviors and expand to domains such as vision and reinforcement learning. Mechanistic interpretability could help prevent catastrophic outcomes as AI systems become more powerful and inscrutable.
@@ 1 Introduction
As AI systems become increasingly sophisticated and general (Bubeck et al., 2023), advancing our understanding of these systems is crucial to ensure their alignment with human values and avoid catastrophic outcomes. The field of interpretability aims to demystify the internal processes of AI models, moving beyond evaluating performance alone. This review focuses on mechanistic interpretability, an emerging approach within the broader interpretability landscape that strives to specify the computations underlying deep neural networks comprehensively. We emphasize that understanding and interpreting these complex systems is not merely an academic endeavor – it’s a societal imperative to ensure AI remains trustworthy and beneficial.
The interpretability landscape is undergoing a paradigm shift akin to the evolution from behaviorism to cognitive neuroscience in psychology. Historically, lacking tools for introspection, psychology treated the mind as a black box, focusing solely on observable behaviors. Similarly, interpretability has predominantly relied on black-box techniques, analyzing models based on input-output relationships or using attribution methods that, while probing deeper, still neglect the model’s internal architecture. However, just as advancements in neuroscience allowed for a deeper understanding of internal cognitive processes, the field of interpretability is now moving towards a more granular approach. This shift from surface-level analysis to a focus on the internal mechanics of deep neural networks characterizes the transition towards inner interpretability (Räuker et al., 2023).
Mechanistic interpretability, as an approach to inner interpretability, aims to completely specify a neural network’s computation, potentially in a format as explicit as pseudocode (also called reverse engineering), striving for a granular and precise understanding of model behavior. It distinguishes itself primarily through its ambition for comprehensive reverse engineering and its strong motivation towards AI safety. Our review serves as the first comprehensive exploration of mechanistic interpretability research, with the most accessible introductions currently scattered in a blog or list format (Olah, 2022; Nanda, 2022d; Olah et al., 2020; Sharkey et al., 2022a; Olah et al., 2018; Nanda, 2023f). We aim to synthesize the research (addressing the "research debt" (Olah & Carter, 2017)) and provide a structured, accessible introduction for AI researchers and practitioners.
@@ 2 Interpretability Paradigms from the Outside In
interpretability aims to explain outputs by tracing predictions to individual input contributions using gradients. Raw gradients can be discontinuous or sensitive to slight perturbations. Therefore, techniques such as SmoothGrad (Smilkov et al., 2017) and Integrated Gradients (Sundararajan et al., 2017) average across gradients. Other popular techniques are layer-wise relevance propagation (Bach et al., 2015), DeepLIFT (Shrikumar et al., 2017), or GradCAM (Selvaraju et al., 2016). Attribution enhances transparency by showing input feature influence without requiring an understanding of the internal structure, enabling decision validation, compliance, and trust while serving as a bias detection tool.
@@ 3.1 Defining Features as Representational Primitives
[htb] Definition 1: Feature Features are the fundamental units of neural network representations that cannot be further decomposed into simpler independent factors.
@@ 3.2 Nature of Features: From Monosemantic Neurons to Non-Linear Representations
[htb] Hypothesis 1: Superposition Neural networks represent more features than they have neurons by encoding features in overlapping combinations of neurons.
@@
[htb] Toy Model of Superposition A toy model (Elhage et al., 2022b) investigates the hypothesis that neural networks can represent more features than the number of neurons by encoding real-world concepts in a compressed manner. The model considers a high-dimensional vector $\mathbf{x}$, where each element $x_{i}$ corresponds to a feature capturing a real-world concept, represented as a random vector with varying importance determined by a weight $a_{i}$. These features are assumed to have the following properties: 1) Concept sparsity: Real-world concepts occur sparsely. 2) More concepts than neurons: The number of potential concepts vastly exceeds the available neurons. 3) Varying concept importance: Some concepts are more important than others for the task at hand. The input vector $\mathbf{x}$ represents features capturing these concepts, defined by a sparsity level $S$ and an importance level $a_{i}$ for each feature $x_{i}$, reflecting the sparsity and varying importance of the underlying concepts. The model dynamics involve transforming $\mathbf{x}$ into a hidden representation $\mathbf{h}$ of lower dimension, and then reconstructing it as $\mathbf{x^{\prime}}$: $\mathbf{h}=W\mathbf{x},\quad\mathbf{x^{\prime}}=\text{ReLU}(W^{T}\mathbf{h}+\mathbf{b})$ The network’s performance is evaluated using a loss function $\mathcal{L}$ weighted by the feature importances $a_{i}$, reflecting the importance of the underlying concepts: $\mathcal{L}=\sum_{x}\sum_{i}a_{i}(x_{i}-x^{\prime}_{i})^{2}$ This toy model highlights neural networks’ ability to encode numerous features representing real-world concepts into a compressed representation, providing insights into the superposition phenomenon observed in neural networks trained on real data. [Uncaptioned image] Figure 4: Illustration of the toy model architecture and the effects of sparsity. (left) Transformation of a five-feature input vector $\mathbf{x}$ into a two-dimensional hidden representation $\mathbf{h}$, and its reconstruction as $\mathbf{x}^{\prime}$ using the weight matrix $W$ and its transpose, with feature importance indicated by a color gradient from yellow to green. (right) The effect of increasing feature sparsity $S$ on the encoding capacity of the network, highlighting the network’s enhanced ability to represent features in superposition as sparsity increases from 00 to $0.9$, illustrated by arrows in the activation space $\mathbf{h}$, which correspond to the columns of the matrix $W$.
@@
[htb] Definition 2: Feature (Alternative) Features are elements that a network would ideally assign to individual neurons if neuron count were not a limiting factor (Bricken et al., 2023). In other words, features correspond to the disentangled concepts that a larger, sparser network with sufficient capacity would learn to represent with individual neurons.
@@
[htb] Hypothesis 2: Linear Representation Features are directions in activation space, i.e. linear combinations of neurons.
@@
Establishing non-linearity can be challenging. For example, Li et al. (2023a) initially found that in a GPT model trained on Othello, the board state could only be decoded with a non-linear probe when represented in terms of "black" and "white" pieces, seemingly violating the linearity assumption. However, Nanda (2023c); Nanda et al. (2023b) later showed that a linear probe sufficed when the board state was decoded in terms of "own" and "opponent’s" pieces, reaffirming the linear representation hypothesis in this case. In contrast, the work by Engels et al. (2024) provides a clear and convincing existence proof for non-linear, multi-dimensional representations in language models.
Despite these exceptions, empirical evidence largely supports the linear representation hypothesis in many contexts, especially for feedforward networks with ReLU activations. Semantic vector calculus in word embeddings (Mikolov et al., 2013), successful linear probing (Alain & Bengio, 2016; Belinkov, 2021), sparse dictionary learning (Bricken et al., 2023; Cunningham et al., 2024; Deng et al., 2023), and linear decoding of concepts (O’Mahony et al., 2023), tasks (Hendel et al., 2023), functions (Todd et al., 2023), sentiment (Tigges et al., 2023), and relations (Hernandez et al., 2023; Chanin et al., 2023) in large language models all point to the prevalence of linear representations. Moreover, linear addition techniques for model steering (Turner et al., 2023; Sakarvadia et al., 2023a; Li et al., 2023b) and representation engineering (Zou et al., 2023) highlight the practical implications of linear feature representations.
@@ 3.3 Circuits as Computational Primitives and Motifs as Universal Circuit Patterns
[htb] Definition 3: Circuit Circuits are sub-graphs of the network, consisting of features and the weights connecting them.
The decomposition of neural networks into circuits for interpretability has shown significant promise, particularly in small models trained for specific tasks such as addition, as seen in the work of Nanda et al. (2023a) and Quirke & Barez (2023). Scaling such a comprehensive circuit analysis to broader behaviors in large language models remains challenging. However, there has been notable progress in scaling circuit analysis of narrow behaviors to larger circuits and models, such as indirect object identification (Wang et al., 2023) and greater-than computations (Hanna et al., 2023) in GPT-2 and multiple-choice question answering in Chinchilla (Lieberum et al., 2023).
In search of general and universal circuits, researchers focus particularly on more general and transferable behaviors. McDougall et al. (2023)’s work on copy suppression in GPT-2’s attention heads sheds light on model calibration and self-repair mechanisms. Davies et al. (2023) and Feng & Steinhardt (2023) focus on how large language models represent symbolic knowledge through variable binding and entity-attribute binding, respectively. Yu et al. (2023) explore mechanisms for factual recall, revealing how circuits dynamically balance pre-trained knowledge with new contextual information. Lan & Barez (2023) extend circuit analysis to sequence continuation tasks, identifying shared computational structures across semantically related sequences.
@@
[htb] Definition 4: Motif Motifs are repeated patterns within a network, encompassing either features or circuits that emerge across different models and tasks.
@@
[htb] Hypothesis 3: Weak Universality There are underlying principles governing how neural networks learn to solve certain tasks. Models will generally converge on analogous solutions that adhere to the common underlying principles. However, the specific features and circuits that implement these principles can vary across different models based on factors like hyperparameters, random seeds, and architectural choices.
[htb] Hypothesis 4: Strong Universality The same core features and circuits will universally and consistently arise across all neural network models trained on similar tasks and data distributions and using similar techniques, reflecting a set of fundamental computational motifs that neural networks inherently gravitate towards when learning.
@@ 3.4 Emergence of World Models and Simulated Agents
LLMs are sometimes characterized as stochastic parrots (Bender et al., 2021). This label stems from their fundamental operational mechanism of predicting the next word in a sequence, which is seen as relying heavily on memorization. From this viewpoint, LLMs are thought to form complex correlations based on observational data but lack the ability to develop causal models of the world due to their lack of access to interventional data (Pearl, 2009).
@@
An alternative perspective on LLMs comes from the active inference framework (Salvatori et al., 2023), a theory rooted in cognitive science and neuroscience. Active inference postulates that the objective of minimizing prediction error, given enough representative capacity, is adequate for a learning system to develop complex world representations, behaviors, and abstractions. Since language inherently mirrors the world, these models could implicitly construct linguistic and broader world models.
@@
The simulation hypothesis suggests that models designed for prediction, such as LLMs, will eventually simulate the causal processes underlying data creation. Seen as an extension of their drive for efficient compression, this hypothesis implies that adequately trained models like GPT could develop internal world models as a natural outcome of their predictive training (janus, 2022).
@@
[htb] Hypothesis 5: Simulation A model whose objective is text prediction will simulate the causal processes underlying the text creation if optimized sufficiently strongly (janus, 2022).
@@
In addition to theoretical considerations for emergent causal world models (Richens & Everitt, 2024; Nichani et al., 2024), mechanistic interpretability is starting to provide empirical evidence on the types of internal world models that may emerge in LLMs. The ability to internally represent the board state in games like Othello (Li et al., 2023a; Nanda et al., 2023b), create linear abstractions of spatial and temporal data (Gurnee & Tegmark, 2024), and structure complex representations of mazes, demonstrating an understanding of maze topology and pathways (Ivanitskiy et al., 2023) highlight the growing abstraction capabilities of LLMs. Li et al. (2021) identified contextual word representations that function as models of entities and situations evolving throughout a discourse, akin to linguistic models of dynamic semantics. Patel & Pavlick (2022) demonstrated that LLMs can map conceptual domains (e.g., direction, color) to grounded world representations given a few examples, suggesting they learn rich conceptual spaces (Gardenfors, 2004) reflective of the non-linguistic world.
@@
The prediction orthogonality hypothesis further expands on this idea: It posits that prediction-focused models like GPT may simulate agents with various objectives and levels of optimality. In this context, GPT are simulators, simulating entities known as simulacra that can be either agentic or non-agentic, with different objectives from the simulator itself (janus, 2022). The implications of the simulation and prediction orthogonality hypotheses for AI safety and alignment are discussed in Section 6.
@@
[htb] Hypothesis 6: Prediction Orthogonality A model whose objective is prediction can simulate agents who optimize toward any objectives with any degree of optimality (janus, 2022).

### v3

@@ Abstract
Understanding AI systems’ inner workings is critical for ensuring value alignment and safety. This review explores mechanistic interpretability: reverse engineering the computational mechanisms and representations learned by neural networks into human-understandable algorithms and concepts to provide a granular, causal understanding. We establish foundational concepts such as features encoding knowledge within neural activations and hypotheses about their representation and computation. We survey methodologies for causally dissecting model behaviors and assess the relevance of mechanistic interpretability to AI safety. We examine benefits in understanding, control, alignment, and risks such as capability gains and dual-use concerns. We investigate challenges surrounding scalability, automation, and comprehensive interpretation. We advocate for clarifying concepts, setting standards, and scaling techniques to handle complex models and behaviors and expand to domains such as vision and reinforcement learning. Mechanistic interpretability could help prevent catastrophic outcomes as AI systems become more powerful and inscrutable. For an HTML version of the paper, visit https://leonardbereska.github.io/blog/2024/mechinterpreview/.
@@ 1 Introduction
As AI systems rapidly become more sophisticated and general (Bubeck et al., 2023; Bengio et al., 2023), advancing our understanding of these systems is crucial to ensure their alignment (Ji et al., 2024) with human values and avoid catastrophic outcomes (Hendrycks et al., 2023; Hendrycks & Mazeika, 2022). The field of interpretability aims to demystify the internal processes of AI models, moving beyond evaluating performance alone. This review focuses on mechanistic interpretability, an emerging approach within the broader interpretability landscape that strives to comprehensively specify the computations underlying deep neural networks. We emphasize that understanding and interpreting these complex systems is not merely an academic endeavor – it’s a societal imperative to ensure AI remains trustworthy and beneficial.
The interpretability landscape is undergoing a paradigm shift akin to the evolution from behaviorism to cognitive neuroscience in psychology. Historically, lacking tools for introspection, psychology treated the mind as a black box, focusing solely on observable behaviors. Similarly, interpretability has predominantly relied on black-box techniques (Casper et al., 2024), analyzing models based on input-output relationships or using attribution methods that, while probing deeper, still neglect the model’s internal architecture. However, just as advancements in neuroscience allowed for a deeper understanding of internal cognitive processes, the field of interpretability is now moving towards a more granular approach. This shift from surface-level analysis to a focus on the internal mechanics of deep neural networks characterizes the transition towards inner interpretability (Räuker et al., 2023).
Mechanistic interpretability, as an approach to inner interpretability, aims to completely specify a neural network’s computation, potentially in a format as explicit as pseudocode (also called reverse engineering), striving for a granular and precise understanding of model behavior. It distinguishes itself primarily through its ambition for comprehensive reverse engineering and its strong motivation towards AI safety. Our review serves as the first comprehensive exploration of mechanistic interpretability research, with the most accessible introductions currently scattered in a blog or list format (Olah, 2022; Nanda, 2022d; Olah et al., 2020; Sharkey et al., 2022a; Olah et al., 2018; Nanda, 2023f; 2024). Concurrently, Ferrando et al. (2024) and Rai et al. (2024) have also contributed valuable reviews giving concise, technical introductions to mechanistic interpretability in transformer-based language models. Our work complements these efforts by synthesizing the research (addressing the "research debt" (Olah & Carter, 2017)) and providing a structured, accessible, and comprehensive introduction for AI researchers and practitioners.
@@ 2 Interpretability Paradigms from the Outside In
interpretability aims to explain outputs by tracing predictions to individual input contributions using gradients. Raw gradients can be discontinuous or sensitive to slight perturbations. Therefore, techniques such as SmoothGrad (Smilkov et al., 2017) and Integrated Gradients (Sundararajan et al., 2017) average across gradients. Other popular techniques are layer-wise relevance propagation (Bach et al., 2015), DeepLIFT (Shrikumar et al., 2017), or GradCAM (Selvaraju et al., 2016). Attribution enhances transparency by showing input feature influence without requiring an understanding of the internal structure, enabling decision validation, compliance, and trust while serving as a bias detection tool, but also has fundamental limitations (Bilodeau et al., 2024).
@@ 3.1 Defining Features as Representational Primitives
Definition 1: Feature Features are the fundamental units of neural network representations that cannot be further decomposed into simpler independent factors.
@@ 3.2 Nature of Features: From Monosemantic Neurons to Non-Linear Representations
Hypothesis 1: Superposition Neural networks represent more features than they have neurons by encoding features in overlapping combinations of neurons.
@@
Toy Model of Superposition A toy model (Elhage et al., 2022b) investigates the hypothesis that neural networks can represent more features than the number of neurons by encoding real-world concepts in a compressed manner. The model considers a high-dimensional vector $\bm{x}$}, where each element ${x}_{i}$ corresponds to a feature capturing a real-world concept, represented as a random vector with varying importance determined by a weight $a_{i}$. These features are assumed to have the following properties: i.) Concept sparsity: Real-world concepts occur sparsely. ii.) More concepts than neurons: The number of potential concepts vastly exceeds the available neurons. iii.) Varying concept importance: Some concepts are more important than others for the task at hand. The input vector $\bm{x}$} represents features capturing these concepts, defined by a sparsity level $S$ and an importance level $a_{i}$ for each feature ${x}_{i}$, reflecting the sparsity and varying importance of the underlying concepts. The model dynamics involve transforming $\bm{x}$} into a hidden representation $\bm{h}$} of lower dimension, and then reconstructing it as $\bm{x}$$^{\prime}$: ${\bm{h}}={\bm{W}}{\bm{x}},\quad{\bm{x}}^{\prime}=\text{ReLU}({\bm{W}}^{\top}{\bm{h}}+{\bm{b}}).$ The network’s performance is evaluated using a loss function $\mathcal{L}$ weighted by the feature importances $a_{i}$, reflecting the importance of the underlying concepts: $\mathcal{L}=\sum_{{\bm{x}}}\sum_{i}a_{i}({x}_{i}-{x}^{\prime}_{i})^{2}.$ This toy model highlights neural networks’ ability to encode numerous features representing real-world concepts into a compressed representation, providing insights into the superposition phenomenon observed in neural networks trained on real data. [Uncaptioned image] Figure 4: Illustration of the toy model architecture and the effects of sparsity. (left) Transformation of a five-feature input vector $\bm{x}$} into a two-dimensional hidden representation $\bm{h}$}, and its reconstruction as $\bm{x}$$^{\prime}$ using the weight matrix $\bm{W}$} and its transpose, with feature importance indicated by a color gradient from yellow to green. (right) The effect of increasing feature sparsity $S$ on the encoding capacity of the network, highlighting the network’s enhanced ability to represent features in superposition as sparsity increases from 0 to $0.9$, illustrated by arrows in the activation space $\bm{h}$}, which correspond to the columns of the matrix $\bm{W}$}.
@@
Definition 2: Feature (Alternative) Features are elements that a network would ideally assign to individual neurons if neuron count were not a limiting factor (Bricken et al., 2023). In other words, features correspond to the disentangled concepts that a larger, sparser network with sufficient capacity would learn to represent with individual neurons.
@@
Hypothesis 2: Linear Representation Features are directions in activation space, i.e., linear combinations of neurons.
@@
Establishing non-linearity can be challenging. For example, Li et al. (2023a) initially found that in a GPT model trained on Othello, the board state could only be decoded with a non-linear probe when represented in terms of white and black pieces, seemingly violating the linearity assumption. However, Nanda (2023c); Nanda et al. (2023b) later showed that a linear probe sufficed when the board state was decoded in terms of "one’s own" and "the opponent’s" pieces, reaffirming the linear representation hypothesis in this case. In contrast, the work by Engels et al. (2024) provides a clear and convincing existence proof for non-linear, multi-dimensional representations in language models.
@@
Despite these exceptions, empirical evidence largely supports the linear representation hypothesis in many contexts, especially for feedforward networks with ReLU activations. Semantic vector calculus in word embeddings (Mikolov et al., 2013), successful linear probing (Alain & Bengio, 2016; Belinkov, 2021), sparse dictionary learning (Bricken et al., 2023; Cunningham et al., 2024; Deng et al., 2023), and linear decoding of concepts (O’Mahony et al., 2023), tasks (Hendel et al., 2023), functions (Todd et al., 2023), sentiment (Tigges et al., 2024), refusal (Arditi et al., 2024), and relations (Hernandez et al., 2023; Chanin et al., 2023) in large language models all point to the prevalence of linear representations. Moreover, linear addition techniques for model steering (Turner et al., 2023; Sakarvadia et al., 2023a; Li et al., 2023b) and representation engineering (Zou et al., 2023) highlight the practical implications of linear feature representations.
Building upon the linear representation hypothesis, recent work investigated the structural organization of these linear features within activation space. Park et al. (2024) reveal a geometric framework for categorical and hierarchical concepts in large language models. Their findings demonstrate that simple categorical concepts (e.g., mammal, bird) are represented as simplices in the activation space, while hierarchically related concepts are orthogonal. This geometric analysis aligns with earlier observations on feature clustering and splitting in neural networks (Elhage et al., 2022b). It suggests that the linear features are not merely scattered directions but are organized to reflect semantic relationships and hierarchies.
@@ 3.3 Circuits as Computational Primitives and Motifs as Universal Circuit Patterns
Definition 3: Circuit Circuits are sub-graphs of the network, consisting of features and the weights connecting them.
The decomposition of neural networks into circuits for interpretability has shown significant promise, particularly in small models trained for specific tasks such as addition, as seen in the work of Nanda et al. (2023a) and Quirke & Barez (2023). Scaling such a comprehensive circuit analysis to broader behaviors in large language models remains challenging. However, there has been notable progress in scaling circuit analysis of narrow behaviors to larger circuits and models, such as indirect object identification (Wang et al., 2023) and greater-than computations (Hanna et al., 2023) in GPT-2 and multiple-choice question answering (Lieberum et al., 2023).
In search of general and universal circuits, researchers focus particularly on more general and transferable behaviors. McDougall et al. (2023)’s work on copy suppression in GPT-2’s attention heads sheds light on model calibration and self-repair mechanisms. Davies et al. (2023) and Feng & Steinhardt (2023) focus on how large language models represent symbolic knowledge through variable binding and entity-attribute binding, respectively. Yu et al. (2023); Nanda et al. (2023c); Lv et al. (2024); Chughtai et al. (2024); Ortu et al. (2024) explore mechanisms for factual recall, revealing how circuits dynamically balance pre-trained knowledge with new contextual information. Lan & Barez (2023) extend circuit analysis to sequence continuation tasks, identifying shared computational structures across semantically related sequences.
@@
Definition 4: Motif Motifs are repeated patterns within a network, encompassing either features or circuits that emerge across different models and tasks.
@@
Hypothesis 3: Weak Universality There are underlying principles governing how neural networks learn to solve certain tasks. Models will generally converge on analogous solutions that adhere to the common underlying principles. However, the specific features and circuits that implement these principles can vary across different models based on factors like hyperparameters, random seeds, and architectural choices.
Hypothesis 4: Strong Universality The same core features and circuits will universally and consistently arise across all neural network models trained on similar tasks and data distributions and using similar techniques, reflecting a set of fundamental computational motifs that neural networks inherently gravitate towards when learning.
@@ 3.4 Emergence of World Models and Simulated Agents
LLMs are sometimes characterized as stochastic parrots (Bender et al., 2021). This label stems from their fundamental operational mechanism of predicting the next word in a sequence, which is seen as relying heavily on memorization. From this viewpoint, LLMs are thought to form complex correlations based on observational data but cannot develop causal models of the world due to their lack of access to interventional data (Pearl, 2009).
@@
An alternative perspective on LLMs comes from the active inference framework (Salvatori et al., 2023), a theory rooted in cognitive science and neuroscience. Active inference postulates that the objective of minimizing prediction error, given enough representative capacity, is adequate for a learning system to develop complex world representations, behaviors, and abstractions. Since language inherently mirrors the world, these models could implicitly construct linguistic and broader world models (Kulveit et al., 2023).
@@
The simulation hypothesis suggests that models designed for prediction, such as LLMs, will eventually simulate the causal processes underlying data creation. Seen as an extension of their drive for efficient compression, this hypothesis implies that adequately trained models like GPT could develop internal world models as a natural outcome of their predictive training (janus, 2022; Shanahan et al., 2023).
@@
Hypothesis 5: Simulation A model whose objective is text prediction will simulate the causal processes underlying the text creation if optimized sufficiently strongly (janus, 2022).
@@
In addition to theoretical considerations for emergent causal world models (Richens & Everitt, 2024; Nichani et al., 2024), mechanistic interpretability is starting to provide empirical evidence on the types of internal world models that may emerge in LLMs. The ability to internally represent the board state in games like chess (Karvonen, 2024) or Othello (Li et al., 2023a; Nanda et al., 2023b), create linear abstractions of spatial and temporal data (Gurnee & Tegmark, 2024), and structure complex representations of mazes, demonstrating an understanding of maze topology and pathways (Ivanitskiy et al., 2023) highlight the growing abstraction capabilities of LLMs. Li et al. (2021) identified contextual word representations that function as models of entities and situations evolving throughout a discourse, akin to linguistic models of dynamic semantics. Patel & Pavlick (2022) demonstrated that LLMs can map conceptual domains (e.g., direction, color) to grounded world representations given a few examples, suggesting they learn rich conceptual spaces (Gardenfors, 2004) reflective of the non-linguistic world.
@@
The prediction orthogonality hypothesis further expands on this idea: It posits that prediction-focused models like GPT may simulate agents with various objectives and levels of optimality. In this context, GPT are simulators, simulating entities known as simulacra that can be either agentic or non-agentic, with different objectives from the simulator itself (janus, 2022; Shanahan et al., 2023). The implications of the simulation and prediction orthogonality hypotheses for AI safety and alignment are discussed in Section 6.
@@
Hypothesis 6: Prediction Orthogonality A model whose objective is prediction can simulate agents who optimize toward any objectives with any degree of optimality (janus, 2022).